one kg of air is compressed poly tropically(n=1.3)from 1 bar and 27 deg celsius to 3 bar.find work transfer,heat transfer and change in internal energy.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Work Done in Polytropic process = > xRT/(1 – n) = > (1000/32) * 0.0821 * 300/(1 – 1.3) = > -2565.625  Joule.
From
Q = U + W
U = mC(T2 – T1) as Temp. is constant.
U = 0.
Hence 
Q = W = – 2565.25 Joule.